# ICL-Router: In-Context Learned Model Representations for LLM Routing

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional capabilities performance across various tasks, including mathematical reasoning (Liu et al. 2024b; Chervonyi et al. 2025), code generation (DeepSeek-AI et al. 2024; Wang et al. 2025b), logical understanding (Liu et al. 2025a; Wang, Jian, and Yang 2025), and STEM problem-solving (Wu et al. 2025; Ma et al. 2025b). Despite these advances, no single model consistently outperforms others in every domain. Instead, models often exhibit complementary strengths and weaknesses, shaped by differences in their training data, architecture, and optimization.
An emerging line of research explores routing methods. The intuition is straightforward: if models excel in different tasks, then dynamically assigning the most capable model in a model pool to each query might maximize overall performance (Chen et al. 2024b; Lu et al. 2024; Zhuang et al. 2024a; Zhang, Zhan, and Ye 2025). However, in practice, achieving this goal is far from trivial. A key challenge is that effective routing critically depends on an accurate understanding of each model’s capabilities. For example, given two general-purpose models, it is difficult to tell which tasks LLaMA (Grattafiori et al. 2024) excels at versus those where Gemma (Team et al. 2025) performs better, unless both are extensively evaluated across a wide range of tasks. This challenge is further exacerbated by the rapid pace of LLM development. As new models are frequently released from time to time, routing methods must be able to incrementally adapt to these new models. Otherwise, the cost of repeating large-scale evaluations and retraining routers will quickly become prohibitive, ultimately limiting the ability of routing methods to benefit from model scaling.
Although some recent efforts have begun to address this challenge, existing approaches fall short of addressing scalability and evaluation efficiency. For instance, RouterDC (Chen et al. 2024b) uses dual contrastive learning to jointly train query and model embeddings, aligning each query with its optimal model. Similarly, EmbedLLM (Zhuang et al. 2024a) applies binary cross-entropy loss to train an embedding-based router that predicts query–model compatibility. However, both RouterDC and EmbedLLM operate on a fixed set of LLMs; incorporating a new model requires retraining the router. More recently, MODEL-SAT (Zhang, Zhan, and Ye 2025) represents models using a hand-crafted “capability instruction” based on performance on the MMLU benchmark (Hendrycks et al. 2021). This removes the need to retrain the router when adding new models, but requires manual instruction design for other benchmarks and prior knowledge of the capabilities they measure.
@@
To this end, in this paper, we propose ICL-Router, a novel method that leverages in-context learned representations to characterize model capabilities. We hypothesize that a model’s performance on diverse queries can serve as in-context exemplars, as the way a model responds to various queries naturally reflects its unique capability profile. However, including hundreds or thousands of queries directly as in-context information would result in an unmanageably long context window. Inspired by the recent concept of in-context vectors (Zhuang et al. 2024b; Liu et al. 2024a), we instead transform these exemplars into compact in-context vectors, substantially reducing context length. At inference time, the router uses these vectors—compact representations of each model’s capability profile—to estimate the probability that each model can correctly answer a given query. When a new LLM is introduced, it is quickly evaluated on a small set of queries to generate in-context vectors (i.e. its capability profile), which is then fed to the router, enabling immediate inclusion without retraining.
@@ 2 Related Work
In-Context Learning. In-context learning (ICL) enables LLMs to perform new tasks using only demonstrations provided in the prompt. This capability was first effectively demonstrated by GPT-3 (Brown et al. 2020), which achieved strong few-shot and even zero-shot performance across diverse tasks by learning from in-context examples. With the advent of LLMs supporting longer context windows, in-context learning has shown remarkable improvements, especially when provided with hundreds or even thousands of examples (Li et al. 2023a; Chen et al. 2023; Bertsch et al. 2025). Nevertheless, this leads to a substantial increase in both the context window size and the cost of inference. To overcome these limitations, recent research has explored the use of in-context vectors to enable LLMs to process context information more efficiently and flexibly. For example, Zhuang et al. (2024b) propose Vector-ICL, which projects continuous input data from various modalities into the model’s embedding space, allowing LLMs to learn directly from vectorized demonstrations in the prompt. Meanwhile, Liu et al. (2024a) focus on improving context selection and demonstration efficiency, further reducing inference costs while maintaining high performance across a range of in-context learning scenarios. Building on these advances, our work explores the use of in-context vectors as model representations for model routing.
@@
Model Routing. Model routing aims to efficiently allocates queries to the optimal model without invoking all available models. Recent research in this area generally falls into two categories: studies that focus on balancing performance with computational cost (Ong et al. 2024; Feng, Shen, and You 2024; Wang et al. 2025a; Jitkrittum et al. 2025; Zhang, Feng, and You 2025), and those that prioritize maximizing model performance (Jiang, Ren, and Lin 2023; Lu et al. 2024; Chen et al. 2024b; Zhuang et al. 2024a; Zhang, Zhan, and Ye 2025; Zhang et al. 2025c). For the former, GraphRouter (Feng, Shen, and You 2024) utilizes a graph-based approach to allocate queries among models, reducing computational overhead while maintaining performance. MixLLM (Wang et al. 2025a) further optimizes the trade-off between resource consumption and accuracy by dynamically selecting models based on input complexity and cost-effectiveness. Router-R1 (Zhang, Zhan, and Ye 2025) treats routing as a sequential decision process, embodying the router itself as an LLM that alternates between “think” and “route” steps to progressively decompose tasks, invoke models dynamically, and jointly balance performance and cost. For the latter, ZOOTER (Lu et al. 2024) introduces a reward-guided approach to query routing, utilizing tag-based label enhancement to promote greater training stability. RouterDC (Chen et al. 2024b) introduces a dual-contrastive learning approach to better align queries and model representations. EmbedLLM (Zhuang et al. 2024a) proposes an encoder–decoder framework that leverages compact model embeddings and query embeddings to predict routing accuracy. MODEL-SAT (Zhang, Zhan, and Ye 2025) employs aptitude-based instruction tuning to characterize model capabilities and dynamically route instructions to the most suitable LLMs. However, existing performance-oriented routing approaches encounter two major challenges: the model representations they construct are generally overly simplistic, and incorporating new models is inefficient, usually requiring substantial retraining and large-scale inference. To overcome these limitations, the proposed ICL-Router employs a two-stage training process to learn semantically rich model representations and enables the efficient integration of new models using only a small set of representative queries.
@@ 4.1 Datasets
We evaluate our method on 10 benchmarks, including OlympiadBench (Chervonyi et al. 2025), AIME, MBPP (Austin et al. 2021), HumanEval (Zhong et al. 2023), BBH (Suzgun et al. 2022), LogicBench (Parmar et al. 2024), KORBench (Ma et al. 2025a), MMLUPro (Wang et al. 2024), AGIEval (Zhong et al. 2023), and MMLU-CF (Zhao et al. 2024). To evaluate the generalization of ICL-Router, we partition these benchmarks into in-distribution and out-of-distribution (OOD) sets. We designate AIME, HumanEval, KORBench, AGIEval and MMLU-CF as OOD datasets, reserving them exclusively for testing. Referring to the settings in RouterDC (Chen et al. 2024b), the other benchmarks are considered in-distribution and split into training and test sets at a 7:3 ratio. Detailed information about these datasets can be found in Appendix A.1.
@@ 4.2 Implementation Settings
Our framework is composed of three main components: an embedding model, a projector, and an LLM-based router. We adopt Qwen3-Embedding-8B (Zhang et al. 2025d) as the embedding model; notably, this model is used solely to generate query vectors and does not participate in the training process. The projector is implemented as a two-layer multi-layer perceptron (MLP) that facilitates the transformation between the embedding space and the router. For the router, we employ Qwen2.5-7B-Instruct, given its good performance and versatility. For the LLM pool, we consider eight widely-used open-source LLMs.
@@ 4.3 Baselines
  • RouterDC (Chen et al. 2024b): Trains query and model embeddings using dual contrastive learning, pulling queries closer to suitable models and clustering semantically similar queries within the representation space.
  • EmbedLLM (Zhuang et al. 2024a): An encoder–decoder framework that employs compact model and query embeddings to predict model-query compatibility, with the router trained via binary cross-entropy loss.
  • MODEL-SAT (Zhang, Zhan, and Ye 2025): Leverages capability instruction tuning by converting candidate model performance into textual descriptions, which are embedded and passed to a trainable LLM that dynamically predicts the most suitable model for each query.
@@ 4.5 Scalability to New Models
We conduct experiments to assess the scalability of our method on a set of recently released and competitive LLMs, including Falcon-H1-7B-Instruct (Zuo et al. 2025), Gemma3-12B-IT (Team et al. 2025), DeepSeek-R1-Llama-8B (Guo et al. 2025), OpenThinker3-7B (Guha et al. 2025), and AceReason-Nemotron-1.1-7B (Liu et al. 2025b). As illustrated in Figures 2 and 3, ICL-Router demonstrates a consistent upward trajectory in accuracy for both in-distribution and out-of-distribution (OOD) scenarios. For example, in the in-distribution setting, ICL-Router’s accuracy improves from 76.3% to 79.9% as more models are integrated; a similar trend is observed on OOD tasks, where performance rises from 66.4% to 69.9%. In contrast, MODEL-SAT exhibits less stable gains, with accuracy improvements that are inconsistent and sometimes fluctuate as additional models are introduced.
@@ 4.6 Analysis
Embedding Model. The choice of embedding model plays a pivotal role in determining the quality of vector representations used for downstream routing. Specially, we evaluate the impact of various embedding models on both in-distribution and OOD accuracy by comparing five embedding models that differ in parameter size and embedding dimensions, including mxbai-embed-large-v1 (Lee et al. 2024) , bge-m3 (Chen et al. 2024a), stella-en-1.5B-v5 (Zhang et al. 2025a), gte-Qwen2-7B-instruct (Li et al. 2023b), Qwen3-8B-Embedding (Zhang et al. 2025d). As shown in Table 3, a clear trend emerges: as model size increases—from the smallest, mxbai-embed-large-v1 (0.3B/1024), to the largest, Qwen3-8B-Embedding (8B/4096)—in-distribution accuracy improves from 74.45% to 76.30%, while out-of-distribution (OOD) accuracy rises from 64.55% to 66.47%. This stepwise improvement across all five evaluated models indicates that stronger embedding models consistently produce more robust and generalizable vector representations, leading to better downstream routing performance.
@@
Effects of In-Context Exemplar Quantity. To further investigate the impact of the number of in-context exemplars on routing performance, we conduct ablation experiments by varying the quantity of exemplars used to construct each model’s capability profile. Specifically, we assess performance with 100, 300, 500, and 1000 in-context exemplars, analyzing both in-distribution and out-of-distribution (OOD) scenarios. As shown in Figures 4 and 5, increasing the number of in-context exemplars consistently improves both in-distribution and out-of-distribution routing accuracy up to a certain point. For in-distribution tasks, accuracy increases from 74.94% with 100 exemplars to a peak of 76.30% at 500 exemplars, before slightly dropping to 75.75% at 1000 exemplars. A similar trend is observed for out-of-distribution performance, where accuracy rises from 65.36% (100 exemplars) to 66.47% (500 exemplars), but then plateaus or slightly declines to 65.71% with 1000 exemplars. These results suggest that while adding more exemplars generally enriches model capability profiles and enhances routing, there is a point of diminishing returns—likely due to redundancy or increased noise as the exemplar set grows too large. This observation is consistent with findings from recent studies on many-shot in-context learning (Agarwal et al. 2024; Li et al. 2024; Zhang et al. 2025b), which report that including too many context examples can sometimes degrade performance rather than improve it. Overall, using a moderate number of exemplars strikes the best balance between informativeness and efficiency, supporting robust and generalizable routing performance.

### v3

@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional capabilities performance across various tasks, including mathematical reasoning (acemath2024; chervonyi2025goldmedalistperformancesolvingolympiad), code generation (deepseekai2024deepseekcoderv2breakingbarrierclosedsource; wang2025coevolvingllmcoderunit), logical understanding (liu2025logicofthoughtinjectinglogiccontexts; wang2025thought), and STEM problem-solving (wu2025sharp; ma2025generalreasoneradvancingllmreasoning). Despite these advances, no single model consistently outperforms others in every domain. Instead, models often exhibit complementary strengths and weaknesses, shaped by differences in their training data, architecture, and optimization.
An emerging line of research explores routing methods. The intuition is straightforward: if models excel in different tasks, then dynamically assigning the most capable model in a model pool to each query might maximize overall performance (chen2024routerdc; lu-etal-2024-routing; zhuang2024embedllm; zhang2025capability). However, in practice, achieving this goal is far from trivial. A key challenge is that effective routing critically depends on an accurate understanding of each model’s capabilities. For example, given two general-purpose models, it is difficult to tell which tasks LLaMA (grattafiori2024llama) excels at versus those where Gemma (team2025gemma) performs better, unless both are extensively evaluated across a wide range of tasks. This challenge is further exacerbated by the rapid pace of LLM development. As new models are frequently released from time to time, routing methods must be able to incrementally adapt to these new models. Otherwise, the cost of repeating large-scale evaluations and retraining routers will quickly become prohibitive, ultimately limiting the ability of routing methods to benefit from model scaling.
Although some recent efforts have begun to address this challenge, existing approaches fall short of addressing scalability and evaluation efficiency. For instance, RouterDC (chen2024routerdc) uses dual contrastive learning to jointly train query and model embeddings, aligning each query with its optimal model. Similarly, EmbedLLM (zhuang2024embedllm) applies binary cross-entropy loss to train an embedding-based router that predicts query–model compatibility. However, both RouterDC and EmbedLLM operate on a fixed set of LLMs; incorporating a new model requires retraining the router. More recently, MODEL-SAT (zhang2025capability) represents models using a hand-crafted “capability instruction” based on performance on the MMLU benchmark (hendryckstest2021). This removes the need to retrain the router when adding new models, but requires manual instruction design for other benchmarks and prior knowledge of the capabilities they measure.
@@
To this end, in this paper, we propose ICL-Router, a novel method that leverages in-context learned representations to characterize model capabilities. We hypothesize that a model’s performance on diverse queries can serve as in-context exemplars, as the way a model responds to various queries naturally reflects its unique capability profile. However, including hundreds or thousands of queries directly as in-context information would result in an unmanageably long context window. Inspired by the recent concept of in-context vectors (zhuang2024vector; liu2024context), we instead transform these exemplars into compact in-context vectors, substantially reducing context length. At inference time, the router uses these vectors—compact representations of each model’s capability profile—to estimate the probability that each model can correctly answer a given query. When a new LLM is introduced, it is quickly evaluated on a small set of queries to generate in-context vectors (i.e. its capability profile), which is then fed to the router, enabling immediate inclusion without retraining.
@@ 2 Related Work
In-Context Learning. In-context learning (ICL) enables LLMs to perform new tasks using only demonstrations provided in the prompt. This capability was first effectively demonstrated by GPT-3 (brown2020language), which achieved strong few-shot and even zero-shot performance across diverse tasks by learning from in-context examples. With the advent of LLMs supporting longer context windows, in-context learning has shown remarkable improvements, especially when provided with hundreds or even thousands of examples (li2023incontextlearningdemonstrationexamples; chen-etal-2023-many; bertsch2025context). Nevertheless, this leads to a substantial increase in both the context window size and the cost of inference. To overcome these limitations, recent research has explored the use of in-context vectors to enable LLMs to process context information more efficiently and flexibly. For example, zhuang2024vector propose Vector-ICL, which projects continuous input data from various modalities into the model’s embedding space, allowing LLMs to learn directly from vectorized demonstrations in the prompt. Meanwhile, liu2024context focus on improving context selection and demonstration efficiency, further reducing inference costs while maintaining high performance across a range of in-context learning scenarios. Building on these advances, our work explores the use of in-context vectors as model representations for model routing.
@@
Model Routing. Model routing aims to efficiently allocates queries to the optimal model without invoking all available models. Recent research in this area generally falls into two categories: studies that focus on balancing performance with computational cost (ong2024routellm; feng2024graphrouter; wang2025mixllm; jitkrittum2025universal; zhang2025router), and those that prioritize maximizing model performance (llm-blender-2023; lu-etal-2024-routing; chen2024routerdc; zhuang2024embedllm; zhang2025capability; zhang2025avengers). For the former, GraphRouter (feng2024graphrouter) utilizes a graph-based approach to allocate queries among models, reducing computational overhead while maintaining performance. MixLLM (wang2025mixllm) further optimizes the trade-off between resource consumption and accuracy by dynamically selecting models based on input complexity and cost-effectiveness. Router-R1 (zhang2025capability) treats routing as a sequential decision process, embodying the router itself as an LLM that alternates between “think” and “route” steps to progressively decompose tasks, invoke models dynamically, and jointly balance performance and cost. For the latter, ZOOTER (lu-etal-2024-routing) introduces a reward-guided approach to query routing, utilizing tag-based label enhancement to promote greater training stability. RouterDC (chen2024routerdc) introduces a dual-contrastive learning approach to better align queries and model representations. EmbedLLM (zhuang2024embedllm) proposes an encoder–decoder framework that leverages compact model embeddings and query embeddings to predict routing accuracy. MODEL-SAT (zhang2025capability) employs aptitude-based instruction tuning to characterize model capabilities and dynamically route instructions to the most suitable LLMs. However, existing performance-oriented routing approaches encounter two major challenges: the model representations they construct are generally overly simplistic, and incorporating new models is inefficient, usually requiring substantial retraining and large-scale inference. To overcome these limitations, the proposed ICL-Router employs a two-stage training process to learn semantically rich model representations and enables the efficient integration of new models using only a small set of representative queries.
@@ 4.1 Datasets
We evaluate our method on 10 benchmarks, including OlympiadBench (chervonyi2025goldmedalistperformancesolvingolympiad), AIME, MBPP (austin2021programsynthesislargelanguage), HumanEval (zhong2023agievalhumancentricbenchmarkevaluating), BBH (suzgun2022challengingbigbenchtaskschainofthought), LogicBench (parmar2024logicbenchsystematicevaluationlogical), KORBench (ma2025korbenchbenchmarkinglanguagemodels), MMLUPro (wang2024mmluprorobustchallengingmultitask), AGIEval (zhong2023agievalhumancentricbenchmarkevaluating), and MMLU-CF (zhao2024mmlucfcontaminationfreemultitasklanguage). To evaluate the generalization of ICL-Router, we partition these benchmarks into in-distribution and out-of-distribution (OOD) sets. We designate AIME, HumanEval, KORBench, AGIEval and MMLU-CF as OOD datasets, reserving them exclusively for testing. Referring to the settings in RouterDC (chen2024routerdc), the other benchmarks are considered in-distribution and split into training and test sets at a 7:3 ratio. Detailed information about these datasets can be found in Appendix A.1.
@@ 4.2 Implementation Settings
Our framework is composed of three main components: an embedding model, a projector, and an LLM-based router. We adopt Qwen3-Embedding-8B (qwen3embedding) as the embedding model; notably, this model is used solely to generate query vectors and does not participate in the training process. The projector is implemented as a two-layer multi-layer perceptron (MLP) that facilitates the transformation between the embedding space and the router. For the router, we employ Qwen2.5-7B-Instruct, given its good performance and versatility. For the LLM pool, we consider eight widely-used open-source LLMs.
@@ 4.3 Baselines
  • RouterDC (chen2024routerdc): Trains query and model embeddings using dual contrastive learning, pulling queries closer to suitable models and clustering semantically similar queries within the representation space.
  • EmbedLLM (zhuang2024embedllm): An encoder–decoder framework that employs compact model and query embeddings to predict model-query compatibility, with the router trained via binary cross-entropy loss.
  • MODEL-SAT (zhang2025capability): Leverages capability instruction tuning by converting candidate model performance into textual descriptions, which are embedded and passed to a trainable LLM that dynamically predicts the most suitable model for each query.
@@ 4.5 Scalability to New Models
We conduct experiments to assess the scalability of our method on a set of recently released and competitive LLMs, including Falcon-H1-7B-Instruct (falconh1), Gemma3-12B-IT (team2025gemma), DeepSeek-R1-Llama-8B (guo2025deepseek), OpenThinker3-7B (guha2025openthoughtsdatarecipesreasoning), and AceReason-Nemotron-1.1-7B (liu2025acereason). As illustrated in Figures 2 and 3, ICL-Router demonstrates a consistent upward trajectory in accuracy for both in-distribution and out-of-distribution (OOD) scenarios. For example, in the in-distribution setting, ICL-Router’s accuracy improves from 76.3% to 79.9% as more models are integrated; a similar trend is observed on OOD tasks, where performance rises from 66.4% to 69.9%. In contrast, MODEL-SAT exhibits less stable gains, with accuracy improvements that are inconsistent and sometimes fluctuate as additional models are introduced.
@@ 4.6 Analysis
Embedding Model. The choice of embedding model plays a pivotal role in determining the quality of vector representations used for downstream routing. Specially, we evaluate the impact of various embedding models on both in-distribution and OOD accuracy by comparing five embedding models that differ in parameter size and embedding dimensions, including mxbai-embed-large-v1 (emb2024mxbai) , bge-m3 (bge-m3), stella-en-1.5B-v5 (zhang2025jasperstelladistillationsota), gte-Qwen2-7B-instruct (li2023towards), Qwen3-8B-Embedding (qwen3embedding). As shown in Table 3, a clear trend emerges: as model size increases—from the smallest, mxbai-embed-large-v1 (0.3B/1024), to the largest, Qwen3-8B-Embedding (8B/4096)—in-distribution accuracy improves from 74.45% to 76.30%, while out-of-distribution (OOD) accuracy rises from 64.55% to 66.47%. This stepwise improvement across all five evaluated models indicates that stronger embedding models consistently produce more robust and generalizable vector representations, leading to better downstream routing performance.
@@
Effects of In-Context Exemplar Quantity. To further investigate the impact of the number of in-context exemplars on routing performance, we conduct ablation experiments by varying the quantity of exemplars used to construct each model’s capability profile. Specifically, we assess performance with 100, 300, 500, and 1000 in-context exemplars, analyzing both in-distribution and out-of-distribution (OOD) scenarios. As shown in Figures 4 and 5, increasing the number of in-context exemplars consistently improves both in-distribution and out-of-distribution routing accuracy up to a certain point. For in-distribution tasks, accuracy increases from 74.94% with 100 exemplars to a peak of 76.30% at 500 exemplars, before slightly dropping to 75.75% at 1000 exemplars. A similar trend is observed for out-of-distribution performance, where accuracy rises from 65.36% (100 exemplars) to 66.47% (500 exemplars), but then plateaus or slightly declines to 65.71% with 1000 exemplars. These results suggest that while adding more exemplars generally enriches model capability profiles and enhances routing, there is a point of diminishing returns—likely due to redundancy or increased noise as the exemplar set grows too large. This observation is consistent with findings from recent studies on many-shot in-context learning (agarwal2024many; li2024long; zhang2025more), which report that including too many context examples can sometimes degrade performance rather than improve it. Overall, using a moderate number of exemplars strikes the best balance between informativeness and efficiency, supporting robust and generalizable routing performance.